# A phase transition for the two-dimensional random field Ising/FK-Ising model

### v1

@@ 1 Introduction
Despite the significant progress regarding the Imry-Ma predictions, some relevant properties of the RFIM are still not fully comprehended. One such property is the near-critical behavior of the two-dimensional RFIM. More specifically, examining what happens when the disorder strength $\epsilon$ varies with the box size $N$ at different rates. In this direction, [14], [16] and [11] study the correlation length—roughly speaking, when critical order that the boundary influence “perceive” the external field, see Sections 2.2 and 2.6 for detailed definitions—of the two-dimensional RFIM for all $T\leq T_{c}.$
@@ 1.1 The random field Ising model: a warm-up
Let $\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}$ and $\mu^{\xi,0}_{T,\Lambda_{N}}$ stand for the two-dimensional Ising measures (see Section 2.2 for the detailed definition) with and without random external field $\epsilon h$, where $h=\{h_{x}:x\in\mathbb{Z}^{d}\}$ is a family of i.i.d. standard normal variables with law $\mathbb{P}$. Our first result concerns the phase transition of the total variation (denoted by $\|\cdot\|_{\mathrm{TV}}$) distance between them.
@@ 1.2 Main results
$\mathbb{P}\left(\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}>c^{-1}\sqrt{\epsilon N^{\alpha(T)}}\right)\leq c^{-1}\exp\left(-c\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right).$
$\mathbb{P}\left(\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}<1-c^{-1}\sqrt{\epsilon^{-1}N^{-\alpha(T)}}\right)\leq c^{-1}\exp\left(-c\epsilon N^{\alpha(T)}\right).$
@@ 1.3 Organization of the paper
In Section 2, we first present some notations and precise definitions of the RFIM and the RFFKIM, along with the fundamental properties of these models. Subsequently, in Section 3, we will prove Theorem 1.1 as a warm-up, since it reveals the core concepts of deriving upper and lower bounds of total variation distance through relatively straightforward calculations. Section 4 will be dedicated to proving the upper bound of the total variation distance between the FK-Ising model with and without external field. In Section 5, we establish the corresponding lower bounds, utilizing a coarse-graining framework to accumulate small differences on small boxes. Finally, in Section 6, we demonstrate that such differences indeed exist due to the effect of the external field on boxes of an appropriate size. In Appendix A, we prove a new LDP result for the FK-Ising model at critical temperature, which is a new result of independent interest.
Throughout our proof, we have used several estimates on the FK-Ising model and Gaussian concentration inequalities, which may be considered intuitive or even elementary for experts in these fields; however, despite our best efforts, we could not locate precise references for some of these results. For the sake of completeness, we have included proofs of them in the Appendices B and C.
@@ 2.1 Notation
Let $\mathbb{Z}^{d}:=\{u=(u_{1},u_{2},\cdots,u_{d}):u_{1},u_{2},\cdots,u_{d}\in\mathbb{Z}\}$ denote the $d$-dimensional integer lattice. For two different vertices $u,v\in\mathbb{Z}^{d},$ we say $u,v$ are adjacent (which we denote as $u\sim v$) if their $l_{1}$-distance is $1$. For a vertex set $V\subset\mathbb{Z}^{2},$ let $G=(V,E)$ denote the induced subgraph, that is $E=\{e=\{x,y\}:x,y\in V,x\sim y\}.$ Let $diam(G)=\max_{u,v\in G}\{||u-v||_{\infty}\}.$ We use the notation $\partial_{int}G$ to denote its interior boundary and $\partial_{ext}G$ to denote its exterior boundary. That is, $\partial_{int}G=\{u\in G|u\sim v\mbox{~{}for~{}some~{}}v\in G^{c}\}$ and $\partial_{ext}G=\{v\in G^{c}|v\sim u\mbox{~{}for~{}some~{}}u\in G\}$. For any configuration $\omega\in\mathbb{R}^{G}$ and $G^{\prime}\subset G$ we use the notation $\omega|_{G^{\prime}}$ to denote its restriction on $G^{\prime}$, i.e., $\omega|_{G^{\prime}}(u)=\omega_{u}$ for any $u\in G^{\prime}$.
For two (discrete) probability measures $\mu$ and $\nu$ on the same measurable space $(\Xi,\mathcal{F})$, the total variation distance between them is defined as
$\|\mu-\nu\|_{\mathrm{TV}}=\frac{1}{2}\sum_{\omega\in\Xi}|\mu(\omega)-\nu(\omega)|.$
Throughout the paper, we use the notation $f(x)=\ln(\cosh(x)).$
@@ 2.2 The random field Ising model (RFIM)
where $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}$ is the partition function given by
$\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}=\sum_{\sigma\in\Sigma}e^{-\frac{1}{T}H_{T,G}^{\xi,\epsilon h}(\sigma)}.$ (2.3)
Note that $\mu_{T,G}^{\xi,\epsilon h}$ and $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}$ are random variables depending on the external field $\epsilon h$. We will denote by $\langle\cdot\rangle_{T,G}^{\xi,\epsilon h}$ the expectation taken with respect to $\mu_{T,G}^{\xi,\epsilon h}$. By taking $\epsilon=0$ we get the pure Ising model.
@@ 2.3 The random field FK-Ising model (RFFKIM)
For a subgraph $G=(V,E)$ of $\mathbb{Z}^{d}$, an edge configuration $\omega\in\Xi=\{0,1\}^{E}$ and a subset $E^{\prime}\subset E$, we use the notation $\omega|_{E^{\prime}}$ to denote the restriction of $\omega$ on $E^{\prime}$. In addition, if there exists $V^{\prime}\subset V$ such that $E^{\prime}=\{(x,y)\in E:x,y\in V^{\prime}\}$, then we also write $\omega|_{V^{\prime}}$ to denote the restriction of $\omega$ on $E^{\prime}$.
@@ 2.4 The Edward-Sokal coupling
    $\pi|_{\sigma}(\omega)\propto\prod_{e\in E_{2}}\delta(\omega(e),0)\prod_{e\in E_{1}}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\Big{]}.$
    Thus, under $\pi|_{\sigma}$, the edges in $E_{2}$ must be closed and in addition, edges in $E_{1}$ are open with probability $p$ and closed with probability $1-p$ independently. In other words, the conditional measure for $\omega$ can be viewed as a Bernoulli bond percolation within each spin cluster.
@@
    $\pi|_{\omega}(\sigma)\propto\prod_{e\in\{x,y\}\in E_{3}}p\delta(\sigma_{x},\sigma_{y}).$
    Thus, the conditional measure $\pi|_{\omega}$ is uniform among spin configurations where each open cluster in $\omega$ receives the same spin.
@@
$\pi_{h}(\sigma,\omega)=\frac{1}{\mathcal{Z}_{\pi_{h}}}\prod_{e=\{x,y\}\in E}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\delta(\sigma_{x},\sigma_{y})\Big{]}\times\exp\left(-\frac{1}{T}\sum_{x\in V}\epsilon h_{x}\sigma_{x}\right),$
@@
    $\sum_{\sigma\in\Sigma}\pi_{h}(\sigma,\omega)=\dfrac{1}{\mathcal{Z}_{\pi_{h}}}\prod_{e\in E}p^{\omega(e)}(1-p)^{1-\omega(e)}\prod_{j=1}^{\kappa(\omega)}2\cosh(h_{\mathcal{C}_{j}}/T),$
    where $\cosh(x)=\frac{e^{x}+e^{-x}}{2}$. Therefore, conditioned on a fixed $\omega\in\Xi$, each open cluster $\mathcal{C}_{j}$ must have the same spin and this spin is plus with probability $\frac{\exp(h_{\mathcal{C}_{j}}/T)}{2\cosh(h_{\mathcal{C}_{j}}/T)}$.
@@ 2.5 Basic properties of the model(s)
Here we record some of the standard and well-known properties that will be used repeatedly.
@@
    $P_{\Gamma_{2}}(\,\cdot\,|\xi)=P_{\Gamma_{1}}^{\xi|_{\partial\Gamma_{1}}}(\cdot),$
    where on the right-hand side $\xi|_{\partial\Gamma_{1}}$ is the restriction on $\partial_{ext}\Gamma_{1}$ if $P$ is an Ising measure and $\xi|_{\partial\Gamma_{1}}$ is a partition on $\partial_{int}\Gamma_{1}$ if $P$ is an FK-Ising measure.
@@
    $p_{c}^{\diamond}(e^{\diamond})=p_{c}=\dfrac{\sqrt{2}}{1+\sqrt{2}}.$
@@ 2.6 Boundary influence and correlation length
where $\sigma\in\{-1,1\}^{\Lambda_{N}}$ is the spin configuration, $\epsilon h=\{\epsilon h_{v}:v\in\mathbb{Z}^{d}\}$ is the external field with strength $\epsilon>0$. And we will say the long-range order exists for the RFIM if the boundary influence has a positive lower bound as the box size $N$ goes to infinity. The absence of long-range order in two dimensions for any temperature $T$ and any disorder strength $\epsilon>0$ raises an intriguing question: can we still observe a phase transition phenomenon for the 2D RFIM? The answer is positive, the behavior of boundary influence alters when $\epsilon$ decays with $N$ in different rates. More precisely, [11] proves when $T=T_{c}(2)$:
@@
We will refer to inequalities such as (1.1) as “absolute continuity” and (1.2) as “singularity” for the sake of simplicity, although this terminology is not mathematically rigorous since all the measures discussed here are discrete. Notice that the total variation distance is always between 0 and 1, thus we can always assume $\epsilon N$ is small enough while proving (1.1) and large enough in (1.2).
@@ 3.1 Obtaining absolute continuity from decoupling
for short. Then by chaos expansion technique ([8, Theorem 3.14]) for Ising (see Section 4.1 for an alternative proof), we have
@@
$\mathbb{P}\otimes\mu\left(G(h,\sigma)\leq 1-c_{3}\sqrt{\epsilon N}\right)\leq c_{3}\exp\left(-c_{3}\sqrt{\epsilon^{-1}N^{-1}}\right),$
@@ 3.2 Obtaining singularity from independency
Let $\{\mu_{i}:i=1,2,\cdots,n\}$ be a sequence of distributions on measurable spaces $\{(\mathcal{E}_{i},\mathcal{F}_{i}):i=1,2,\cdots,n\}.$ Then for any event $\mathcal{A}_{i}\in\mathcal{F}_{i}$, we define $\hat{\mathcal{A}}_{i}=\mathcal{E}_{1}\otimes\cdots\mathcal{E}_{i-1}\otimes\mathcal{A}_{i}\otimes\mathcal{E}_{i+1}\cdots\otimes\mathcal{E}_{n}$ to be a cylinder set.
@@
$\|\mu-\nu\|_{\mathrm{TV}}\geq 1-c_{1}\exp\left(-c_{2}^{-1}a^{2}n\right).$
@@
Let $\mu_{1},\mu_{2},\cdots,\mu_{n}$ and $\nu_{1},\nu_{2},\cdots,\nu_{n}$ be two sequences of independent random distributions with joint law $\mathbb{P}$. Assume ${\mathbb{E}}\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq a>0$ for all $1\leq i\leq n$ where $\mathbb{E}$ is the expectation w.r.t. to $\mathbb{P}$, and $\mu=\otimes_{i=1}^{n}\mu_{i},~{}\nu=\otimes_{i=1}^{n}\nu_{i}$ be their product measures. Then there exists an absolute constant $c_{1}>0$ such that
$\mathbb{P}\Big{(}\|\mu-\nu\|_{\mathrm{TV}}\geq 1-c_{1}\exp(-c_{1}^{-1}n)\Big{)}\geq 1-c_{1}\exp\left({-}c_{1}^{-1}a^{2}n\right).$
@@
By DMP, the distribution of points in $U_{N}$ are independent of each other. Let $V_{x}$ denote the set of neighbours of $x$, and write $\sigma_{V_{x}}=\sigma_{V}\mid_{V_{x}}$ then we have

### v2

@@ 1 Introduction
Despite the significant progress regarding the Imry-Ma predictions, some relevant properties of the RFIM are still not fully comprehended. One such property is the near-critical behavior of the two-dimensional RFIM. More specifically, examining what happens when the disorder strength $\epsilon$ varies with the box size $N$ at different rates. In this direction, [18], [20] and [15] study the correlation length—roughly speaking, when critical order that the boundary influence “perceive” the external field, see Section 2.6 for detailed definitions—of the two-dimensional RFIM for all $T\leq T_{c}.$ Furthermore, in [8, 9, 10], they study the Ising model with a uniform external field and show the phase transition of boundary influence with respect to the strength of the external field.
@@ 1.1 The random field Ising model: a warm-up
Let $\mu^{\xi,\epsilon h}_{T,\Lambda_{N}}$ and $\mu^{\xi,0}_{T,\Lambda_{N}}$ stand for the two-dimensional Ising measures (see Section 2.2 for detailed definitions) with and without random external field $\epsilon h$, where $h=\{h_{x}:x\in\mathbb{Z}^{d}\}$ is a family of i.i.d. standard normal variables with law $\mathbb{P}$. Our first result concerns the phase transition of the total variation (denoted by $\|\cdot\|_{\mathrm{TV}}$) distance between them.
@@ 1.2 Main results
$\mathbb{P}\left(\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}>c^{-1}\sqrt{\epsilon N^{\alpha(T)}}\right)\leq c^{-1}\exp\left(-c\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right).$ (1.3)
$\mathbb{P}\left(\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}\|_{\mathrm{TV}}<1-c^{-1}\sqrt{\epsilon^{-1}N^{-\alpha(T)}}\right)\leq c^{-1}\exp\left(-c\epsilon N^{\alpha(T)}\right).$ (1.4)
@@ 1.3 Organization of the paper
In Section 2, we first present some notations and precise definitions of the RFIM and the RFFKIM, along with the fundamental properties of these models. Subsequently, in Section 3, we will prove Theorem 1.1 as a warm-up, since it reveals the core concepts of deriving upper and lower bounds of total variation distance through relatively straightforward calculations. Section 4 will be dedicated to proving the upper bound of the total variation distance between the FK-Ising model with and without external field (i.e. (1.3)). The corresponding lower bounds (i.e. (1.4)) will be postponed to Section 6, utilizing a coarse-graining framework to accumulate small differences on small boxes. Before that, we demonstrate that the differences in small boxes indeed exist due to the effect of the external field on boxes of an appropriate size in Section 5. In Appendix A, we prove an LDP result for the (discrete) FK-Ising model at critical temperature. While similar results exist for the continuous scaling limit [11, Theorem 3], the discrete case cannot be derived directly from it.
Throughout our proof, we have used several estimates on the FK-Ising model and Gaussian concentration inequalities, which may be considered intuitive or even elementary for experts in these fields; however, despite our best efforts, we could not locate precise references for these results. For the sake of completeness, we have included proofs of them in the Appendices B and C.
@@ 2.1 Notation
[…]
For two (discrete) probability measures $\mu$ and $\nu$ on the same measurable space $(\Xi,\mathcal{F})$, the total variation distance between them is defined as $\|\mu-\nu\|_{\mathrm{TV}}=\frac{1}{2}\sum_{\omega\in\Xi}|\mu(\omega)-\nu(\omega)|.$
For simplicity, we denote $f(x)=\ln(\cosh(x))$ throughout the paper, where $\cosh(x)=\frac{e^{x}+e^{-x}}{2}$, as this function appears frequently.
@@ 2.2 The random field Ising model (RFIM)
where $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}$ is the partition function given by $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}=\sum_{\sigma\in\Sigma}e^{-\frac{1}{T}H_{T,G}^{\xi,\epsilon h}(\sigma)}.$ Note that $\mu_{T,G}^{\xi,\epsilon h}$ and $\mathcal{Z}_{\mu,T,G}^{\xi,\epsilon h}$ are random variables depending on the external field $\epsilon h$. We will denote by $\langle\cdot\rangle_{T,G}^{\xi,\epsilon h}$ the expectation taken with respect to $\mu_{T,G}^{\xi,\epsilon h}$. By taking $\epsilon=0$ we get the pure Ising model.
@@ 2.3 The random field FK-Ising model (RFFKIM)
For a subgraph $G=(V,E)$ of $\mathbb{Z}^{d}$, an edge configuration $\omega\in\Xi=\{0,1\}^{E}$ and a subset $E^{\prime}\subset E$, recall that $\omega_{|{E^{\prime}}}$ is the restriction of $\omega$ on $E^{\prime}$. In addition, if there exists $V^{\prime}\subset V$ such that $E^{\prime}=\{(x,y)\in E:x,y\in V^{\prime}\}$, then we also write $\omega_{|V^{\prime}}$ to denote the restriction of $\omega$ on $E^{\prime}$.
@@ 2.4 The Edward-Sokal coupling
    $\pi_{|\sigma}(\omega)\propto\prod_{e\in E_{2}}\delta(\omega(e),0)\prod_{e\in E_{1}}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\Big{]}.$
    Thus, under $\pi_{|\sigma}$, the edges in $E_{2}$ must be closed and in addition, edges in $E_{1}$ are open with probability $p$ and closed with probability $1-p$ independently. In other words, the conditional measure for $\omega$ can be viewed as a Bernoulli bond percolation within each spin cluster.
@@
    $\pi_{|\omega}(\sigma)\propto\prod_{e\in\{x,y\}\in E_{3}}p\delta(\sigma_{x},\sigma_{y}).$
    Thus, the conditional measure $\pi_{|\omega}$ is uniform among spin configurations where each open cluster in $\omega$ receives the same spin.
@@
$\pi_{h}(\sigma,\omega)=\frac{1}{\mathcal{Z}_{\pi_{h}}}\prod_{e=\{x,y\}\in E}\Big{[}(1-p)\delta(\omega(e),0)+p\delta(\omega(e),1)\delta(\sigma_{x},\sigma_{y})\Big{]}\cdot\exp\left(-\frac{1}{T}\sum_{x\in V}\epsilon h_{x}\sigma_{x}\right),$
@@
    $\sum_{\sigma\in\Sigma}\pi_{h}(\sigma,\omega)=\dfrac{1}{\mathcal{Z}_{\pi_{h}}}\prod_{e\in E}p^{\omega(e)}(1-p)^{1-\omega(e)}\prod_{j=1}^{\kappa(\omega)}2\cosh(h_{\mathcal{C}_{j}}/T).$
    Therefore, conditioned on a fixed $\omega\in\Xi$, each open cluster $\mathcal{C}_{j}$ must have the same spin and this spin is plus with probability $\frac{\exp(h_{\mathcal{C}_{j}}/T)}{2\cosh(h_{\mathcal{C}_{j}}/T)}$.
@@ 2.5 Basic properties of the model(s)
Here we state some of the standard and well-known properties that will be used repeatedly.
@@
    $P_{\Gamma_{2}}(\,\cdot\,|\xi)=P_{\Gamma_{1}}^{\xi_{|\partial\Gamma_{1}}}(\cdot),$
    where on the right-hand side $\xi_{|\partial\Gamma_{1}}$ is the restriction on $\partial_{ext}\Gamma_{1}$ if $P$ is an Ising measure and $\xi_{|\partial\Gamma_{1}}$ is a partition on $\partial_{int}\Gamma_{1}$ if $P$ is an FK-Ising measure.
@@
    $p_{c}^{\diamond}=p_{c}=\dfrac{\sqrt{2}}{1+\sqrt{2}}.$
@@ 2.6 Boundary influence and correlation length
where $\sigma\in\{-1,1\}^{\Lambda_{N}}$ is the spin configuration, $\epsilon h=\{\epsilon h_{v}:v\in\mathbb{Z}^{d}\}$ is the external field with strength $\epsilon>0$. And we will say the long-range order exists for the RFIM if the boundary influence has a positive lower bound as the box size $N$ goes to infinity. The absence of long-range order in two dimensions for any temperature $T$ and any disorder strength $\epsilon>0$ raises an intriguing question: can we still observe a phase transition phenomenon for the 2D RFIM? The answer is positive, the behavior of boundary influence alters when $\epsilon$ decays with $N$ in different rates.
More precisely, Define the correlation length of two-dimensional RFIM as follows:
${\psi_{\star}(T,\epsilon)}=\min\{N:m(T,N,\epsilon)\leq m(T,N,0)/2\}\,.$
By a sequence of works [18, 20, 15], they prove that the correlation length has order $ce^{c\epsilon^{-4/3}}$ at low temperature $T<T_{c}$ and has order $c\epsilon^{-8/7}$ at critical temperature $T=T_{c}$. Furthermore, [15] proves when $T=T_{c}(2)$:
@@
We will refer to inequalities such as (1.1) as “absolute continuity” and (1.2) as “singularity” for the sake of simplicity, although this terminology is not mathematically rigorous since all the measures discussed here are discrete. Notice that the total variation distance is always between 0 and 1, thus we can always assume $\epsilon N$ is small enough while proving (1.1) and large enough while proving (1.2).
@@ 3.1 Obtaining absolute continuity from decoupling
for short. Then by chaos expansion technique ([12, Theorem 3.14]) for Ising (see Section 4.1 for a detailed proof), we have
@@
$\mathbb{P}\otimes\mu^{{\xi,0}}_{{T,\Lambda_{N}}}\left(G(h,\sigma)\leq 1-c_{3}\sqrt{\epsilon N}\right)\leq c_{3}\exp\left(-c_{3}\sqrt{\epsilon^{-1}N^{-1}}\right),$
@@ 3.2 Obtaining singularity from independency
Let $\{\mu_{i}:i=1,2,\cdots,n\}$ be a sequence of distributions on measurable spaces $\{(\mathcal{E}_{i},\mathcal{F}_{i}):i=1,2,\cdots,n\}.$ Then for any event $\mathcal{A}_{i}\in\mathcal{F}_{i}$, we define $\hat{\mathcal{A}}_{i}=\mathcal{E}_{1}\otimes\cdots\otimes\mathcal{E}_{i-1}\otimes\mathcal{A}_{i}\otimes\mathcal{E}_{i+1}\otimes\cdots\otimes\mathcal{E}_{n}$ to be a cylinder set.
@@
$\|\mu-\nu\|_{\mathrm{TV}}\geq 1-c_{1}\exp\left(-{c_{1}}^{-1}a^{2}n\right).$
@@
Let $\mu_{1},\mu_{2},\cdots,\mu_{n}$ and $\nu_{1},\nu_{2},\cdots,\nu_{n}$ be two sequences of independent random distributions with joint law $\mathbb{P}$. Assume ${\mathbb{E}}\|\mu_{i}-\nu_{i}\|_{\mathrm{TV}}\geq a>0$ for all $1\leq i\leq n$ where $\mathbb{E}$ is the expectation w.r.t. $\mathbb{P}$, and $\mu=\otimes_{i=1}^{n}\mu_{i},~{}\nu=\otimes_{i=1}^{n}\nu_{i}$ are product measures. Then there exists an absolute constant $c_{1}>0$ such that
$\mathbb{P}\Big{(}\|\mu-\nu\|_{\mathrm{TV}}\geq 1-c_{1}\exp(-c_{1}^{-1}{a^{2}}n)\Big{)}\geq 1-c_{1}\exp\left({-}c_{1}^{-1}a^{2}n\right).$
@@
By DMP, the distribution of points in $U_{N}$ are independent of each other. Let $V_{x}$ denote the set of neighbours of $x$, and write $\sigma_{V_{x}}=(\sigma_{V})_{\mid V_{x}}$ then we have